OpenStudy (anonymous):

solve the system of equations. y = 3x + 2 x - 2y = 10 I hope someone maybe able to help me I am really confused

OpenStudy (welshfella):

first write them in the same form y = 3x + 2 2y = x - 10 now multiply the first equation by -2 then add it to second equation this will eliminate y and you can solve for x

OpenStudy (welshfella):

so we get -2y = -6x - 4 2y = x - 10 adding them eliminate y;- 0 = -5x - 14

OpenStudy (welshfella):

can you solve this for x?

OpenStudy (anonymous):

I am still a little confused sorry

OpenStudy (welshfella):

we are using the elimination method if we can get an equation with the value of term in x or y such that when we add the 2 equations we can eliminate the term is x or y then we get an equation in only one unknown .

OpenStudy (welshfella):

and getting the 2 equations in same form first mad it easier to add them

OpenStudy (anonymous):

I understand a little better now

OpenStudy (welshfella):

ok so we solve 0 = -5x - 14 add 5x to both sides 5x = -14 x = -14/5 = -2.8

OpenStudy (welshfella):

so we have found the value of x to find the value of y we just substitute x = -2.8 into one of the original equations y = 3x + 2 so y = 3(-2.8) + 2

OpenStudy (welshfella):

can you do that calculation?

OpenStudy (anonymous):

-6.4 would be answer

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

so the solution set is {-2.8,-6.4}

OpenStudy (welshfella):

x is first value , y the second